# Boundary-value problems of functional differential equations with state-dependent delays

### v2

@@ 7 Implementation and numerical test
and $M\in\mathbb{R}^{n_{y}\times n_{y}}$, $y(t)\in\mathbb{R}^{n_{y}}$, and $f:\mathbb{R}^{n_{y}\times(n_{\mathrm{d}}+1)}\times\mathbb{R}^{n_{p}}\to\mathbb{R}^{n_{y}}$, $\tau_{\mathrm{fun},j}:\mathbb{R}^{n_{y}\times j}\times\mathbb{R}^{n_{p}}\to\mathbb{R}$ are smooth functions of their arguments. The matrix $M$ may be singular to permit the implicit definition of delays, or the formulation of neutral FDEs (not analyzed in this paper).
For this class of FDEs the abstract right-hand side $G_{\mathrm{FDE}}$ in (1.1) has the form
@@
Figure 7.1: Left: maximum among the errors computed at $10001$ equidistant points in $[0,T]$ for different values of $m$ and $L$, compared to straight dashed lines having slope $m$. Right: periodic solution computed with $L=10$ and $m=5$, rescaled so as to have period 1.
@@ 8 Conclusions
A convergence analysis was recently performed for FDEs with constant delays (Andò and Breda, 2020), based on the general approach for BVPs established in (Maset, 2016). The latter assumes regularity properties on the part of the right-hand sides that cannot be satisfied when state-dependent delays are present. However, using the concepts of weaker forms of differentiability developed for state-dependent delays (Cassidy et al., 2019; Hartung et al., 2006), allowed us to arrive at convergence estimates as sharp as those previously obtained for constant delays by only assuming to satisfy the regularity properties in the milder form. In particular, our main result is the error estimate $O(L^{-\min\{\ell_{\max},m\}})$ for right-hand sides which are mildly differentiable to order $\ell_{\max}$, where $m$ is the fixed degree of the approximating piecewise polynomials and $L$ is the increasing size of the mesh.
@@ Acknowledgments
For $k=0$ the fact that $D^{\ell}{G_{\mathrm{ts}}}|_{C^{\ell}_{\pi}}$ is well defined at each time $t$ follows from the assumption that $G|_{C^{\ell}_{\pi}}$ is $\ell$ times continuously differentiable by part 1 of Definition 2.1. The continuity with respect to the $C^{0}_{\pi}$-norm in $t$ follows from the fact that the time shift ×$\mathbb{R}\times C^{j}_{\pi}\ni(t,u)\mapsto u_{t}\in C^{j}_{\pi}$ is continuous for arbitrary $j\geq 0$.
@@
By Lemma A.1 the derivative $D^{\ell+1}G(u)v_{0}\ldots v_{j}$ is continuous for arguments $u\in C^{\ell+1}_{\pi}$, $v_{0},\ldots,v_{\ell}\in C^{\ell}_{\pi}$. The argument $u=sx_{t+h}+(1-s)x_{t}$ is in $C^{k+\ell+1}_{\pi}$, $v_{0}=[x_{t+h}-x_{t}]/h$ is in $C^{k+\ell+1}_{\pi}$ and $y^{i}_{t}$ and the limit $x_{t}^{\prime}$ of $[x_{t+h}-x_{t}]/h$ are in $C^{k+\ell}_{\pi}$ such that we can take the limit for $h\to 0$ to obtain
@@
In the second sum in (A.1) the difference quotient $[y^{j}_{t+h}-y^{j}_{t}]/h$ has the limit $(y^{j}_{t})^{\prime}$ in $C^{k+\ell-1}_{\pi}$ and the derivative $D^{\ell}G(x_{t})$ can be extended continuously to multilinear arguments in $C^{k+\ell-1}_{\pi}$, since $G$ is $k+\ell\geq\ell$ times mildly differentiable. Thus, the limit for $h\to 0$ also exists for the second sum in (A.1):
@@
The argument $x$ in this limit is in $C^{k+\ell+1}_{\pi}$, the arguments $y^{i}_{t}$ are in $C^{k+\ell}_{\pi}$ and $(y^{j})^{\prime}$ is in $C^{k+\ell-1}_{\pi}$. Thus, by the assumption in the inductive step the limits in (A.2) are also in $C^{k}_{\pi}$ for $j=1\ldots,\ell$. This implies that the limits of the both terms in (A.1), and, thus, the limit of the time difference quotient (A.1) are in $C^{k}_{\pi}$. Hence, $t\mapsto D^{\ell}G(x_{t})y^{1}_{t}\ldots y^{\ell}_{t}$ is in $C^{k+1}_{\pi}$. Finally, we check the continuity of the map $C^{k+\ell+1}_{\pi}\times(C^{k+\ell}_{\pi})^{\ell}\ni(x,y^{1},\ldots,y^{\ell})\mapsto D^{\ell}{G_{\mathrm{ts}}}(x)y^{1}\ldots y^{\ell}\in C^{k+1}_{\pi}$. By inductive assumption the map is continuous as a map into $C^{k}_{\pi}$. Inspecting (A.1) and (A.2), its highest derivative in $t$ is

### v3

@@ 7 Implementation and numerical test
and $M\in\mathbb{R}^{n_{y}\times n_{y}}$, $y(t)\in\mathbb{R}^{n_{y}}$, and $f:\mathbb{R}^{n_{y}\times(n_{\mathrm{d}}+1)}\times\mathbb{R}^{n_{p}}\to\mathbb{R}^{n_{y}}$, $\tau_{\mathrm{fun},j}:\mathbb{R}^{n_{y}\times j}\times\mathbb{R}^{n_{p}}\to\mathbb{R}$ are smooth functions of their arguments. The matrix $M$ may be singular to permit the implicit definition of delays, or the formulation of neutral FDEs (not analyzed in this paper). For this class of FDEs the abstract right-hand side $G_{\mathrm{FDE}}$ in (1.1) has the form
@@
Figure 7.1: Left: maximum among the errors computed at 1000110001 equidistant points in $[0,T]$ for different values of $m$ and $L$, compared to straight dashed lines having slope $m$. Right: periodic solution computed with $L=10$ and $m=5$, rescaled so as to have period 1.
@@ 8 Conclusions
A convergence analysis was recently performed for FDEs with constant delays (Andò and Breda, 2020), based on the general approach for BVPs established in (Maset, 2016). The latter assumes regularity properties on the part of the right-hand sides that cannot be satisfied when state-dependent delays are present. However, using the concepts of weaker forms of differentiability developed for state-dependent delays (Cassidy et al., 2019; Hartung et al., 2006), allowed us to arrive at convergence estimates as sharp as those previously obtained for constant delays by only assuming to satisfy the regularity properties in the milder form.
In numerical bifurcation analysis the next step after finding a periodic orbit numerically, is consideration of its linear stability, which ensures local asymptotic stability with respect to compatible perturbations in the initial value (Mallet-Paret and Nussbaum, 2011). For numerical computation of linear stability one needs to approximate the monodromy operator and the Floquet multipliers. Although we are not aware of any theoretical study on the convergence of such approximation with respect to the FEM strategy, preliminary experimental evidence can be found in (Breda et al., 2022).
@@ Acknowledgments
For $k=0$ the fact that $D^{\ell}{G_{\mathrm{ts}}}|_{C^{\ell}_{\pi}}$ is well defined at each time $t$ follows from the assumption that $G|_{C^{\ell}_{\pi}}$ is $\ell$ times continuously differentiable by part 1 of Definition 2.1. The continuity with respect to the $C^{0}_{\pi}$-norm in $t$ follows from the fact that the time shift $\mathbb{R}\times C^{j}_{\pi}\ni(t,u)\mapsto u_{t}\in C^{j}_{\pi}$ is continuous for arbitrary $j\geq 0$.
@@
By Lemma A.1 the derivative $D^{\ell+1}G(u)v_{0}\ldots v_{j}$ is continuous for arguments $u\in C^{\ell+1}_{\pi}$, $v_{0},\ldots,v_{\ell}\in C^{\ell}_{\pi}$. The argument $u=sx_{t+h}+(1-s)x_{t}$ is in $C^{k+\ell+1}_{\pi}$, $v_{0}=[x_{t+h}-x_{t}]/h$ is in $C^{k+\ell+1}_{\pi}$ and $y^{i}_{t}$ and the limit $x_{t}^{\prime}$ of $x_{t+h}-x_{t}]/h$ are in $C^{k+\ell}_{\pi}$ such that we can take the limit for $h\to 0$ to obtain
@@
In the second sum in (A.1) the difference quotient $y^{j}_{t+h}-y^{j}_{t}]/h$ has the limit $(y^{j}_{t})^{\prime}$ in $C^{k+\ell-1}_{\pi}$ and the derivative $D^{\ell}G(x_{t})$ can be extended continuously to multilinear arguments in $C^{k+\ell-1}_{\pi}$, since $G$ is $k+\ell\geq\ell$ times mildly differentiable. Thus, the limit for $h\to 0$ also exists for the second sum in (A.1):
@@
The argument $x$ in this limit is in $C^{k+\ell+1}_{\pi}$, the arguments $y^{i}_{t}$ are in $C^{k+\ell}_{\pi}$ and $(y^{j})^{\prime}$ is in $C^{k+\ell-1}_{\pi}$. Thus, by the assumption in the inductive step the limits in (A.2) are also in $C^{k}_{\pi}$ for $j=1\ldots,\ell$. This implies that the limits of the both terms in (A.1), and, thus, the limit of the time difference quotient (A.1) are in $C^{k}_{\pi}$. Hence, $t\mapsto D^{\ell}G(x_{t})y^{1}_{t}\ldots y^{\ell}_{t}$ is in $C^{k+1}_{\pi}$. Finally, we check the continuity of the map $C^{k+\ell+1}_{\pi}\times(C^{k+\ell}_{\pi})^{\ell}\ni(x,y^{1},\ldots,y^{\ell})\mapsto D^{\ell}{G_{\mathrm{ts}}}(x)y^{1}\ldots y^{\ell}\in C^{k+1}_{\pi}$. By inductive assumption the map is continuous as a map into $C^{k}_{\pi}$. Inspecting (A.1) and (A.2), its derivative in $t$ is